OpenStudy (anonymous):

Select the equations that are parallel and perpendicular to y = -3x - 1 and that pass through the point (3, 1). A) parallel: y = -3x + 6 perpendicular: y =1/3x+2 2/3 B)parallel: y = -3x + 10 perpendicular: y =1/3x C) parallel: y =-1/3x perpendicular: y = -3x D)parallel: y =1/3x+1 perpendicular: y = 3x - 1

OpenStudy (campbell_st):

parallel lines have the same slope you line has a slope of m = -3 the perpendicular line is the negative reciprocal... so if m = -3 then the perpendicular line has a slope of m = 1/3 the point slope form of a line is y = mx + b you should look for a parallel line in the form y = -3x + b and a perpendicular line in the form y = 1/3x + c

OpenStudy (campbell_st):

to find the values of b and c substitute x = 3 and y = 1 into both equations